OpenStudy (anonymous):

cube root 24x^8 Someone please help me :(

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

man i messed that all up. lets do it the easy way

OpenStudy (anonymous):

\[24=8\times 3\] and \[\sqrt[3]{8}=2\] so a 2 comes out of the radical and a 3 stays in

OpenStudy (anonymous):

lol ok easy sounds good

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

3 goes in to 8 two times with a remainder of 2, so x^2 comes out and an x^2 stays in

OpenStudy (anonymous):

the answer is \[2x^2\sqrt[3]{3x^2}\]

OpenStudy (anonymous):

i bet this is confusing you

OpenStudy (anonymous):

not really i get it i just wanted to make sure I wasnt doing it wrong thanks so much you are awesome!

OpenStudy (anonymous):

yw sorry i screwed up at first

OpenStudy (anonymous):

do u think you can help me with the other question I had?

OpenStudy (anonymous):

\[\sqrt[3]{24x^8}\], now see here . \[\sqrt[3]{24x^8}\] now we have \[x^8=(x^2)^3\] \[\sqrt[3]{24x^8}\] that is : \[\sqrt[3]{3*2*2*2*(x^2)^3}\] we have now \[x^2*2\sqrt[3]{3}\] hence the answer is \[x^2*2\sqrt[3]{3}\] got it

OpenStudy (anonymous):

u can ask more questions miizzt